OpenStudy (zaarin):

A floor is covered by 600 tiles which are 10 cm by 10 cm. How many 20 cm by 20 cm tiles are needed to cover the same floor? (use similarity formula)

OpenStudy (anonymous):

\[area of floor=10(10)(600)=60000\] \[20(20)x=60000\] \[x=\frac{ 60000 }{ 20(20) }=150\]

OpenStudy (anonymous):

10 x 10 cm = 100 sq, cm 20 x 20 cm = 400 sq. cm. So you need 1/4 of # of tiles..or 150 tiles.

OpenStudy (anonymous):

Find the area of the floor, area of 10 by 10 is 100 and since there are 600 tiles, the area of the floor is 60000. Now find the area of a 20 by 20 tile which is 400 and find the amount of tiles necessary to cover 60000. 60000/400 = 600/4 = 150 tiles

OpenStudy (zaarin):

Thank you so much everyone

OpenStudy (anonymous):

My pleasure :)